OpenStudy (anonymous):

Find the fourth term of the geometric sequence whose first term is 7 and whose common ratio is -3

OpenStudy (ash2326):

A GP Is in the form of a, ar, ar^2.... a= first term r= common ratio its nth term is given as \[a_n=ar^{n-1}\] We are given first term as a=7 r=-3 We have to find 4th term \[a_4=7\times (-3)^{(4-1)}\] \[a_4=7 \times (-3)^3=7\times (-27)=-189\] so the fourth term is -189